OpenStudy (anonymous):

Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura keeps the movie for 3 nights, how much will she pay when she returns it? $2.50 $6.50 $ 3.50 $7.50

OpenStudy (anonymous):

This is almost like your last one, but easier :) You can just add up how much she spends, night by night.

OpenStudy (anonymous):

If she keeps it for zero nights, she still has to pay the flat fee of $2, right? If she keeps it for 1 night, she pays the flat fee of $2 plus the nightly fee of $0.50, or $2.50 total. It's the same idea for nights 2 and 3.

OpenStudy (anonymous):

If you want to see EXACTLY how it's like the last one, the cost function here would be c(x) = 0.50 x + 2 <<-- 0.50 per night, plus a flat $2 Then for 3 nights, x = 3, so cost is c(3) = 0.50(3) + 2 = 1.50 + 2 = 3.50